OpenStudy (anonymous):

(a) Two point charges totaling 7.00 µC exert a repulsive force of 0.200 N on one another when separated by 0.215 m. What is the charge on each? smallest charge Incorrect: Your answer is incorrect. µC largest charge Incorrect: Your answer is incorrect. µC (b) What is the charge on each if the force is attractive? smallest charge µC largest charge µC

OpenStudy (anonymous):

suppose the charges are q1 and q2 the question tells you the value of q1 + q2 also, knowing the force between them, you can use coulombs law to calculate q1*q2 you then have 2 equations for your 2 unknowns, q1 and q2

OpenStudy (anonymous):

can you show me the formula revised to figure it out

OpenStudy (anonymous):

do you know coulombs law for the force between two charges ?

OpenStudy (anonymous):

i know f=kq1g2/r^2

OpenStudy (anonymous):

then its a quadratic right?

OpenStudy (anonymous):

yes, you do end up with a quadratic for the values of the charge

OpenStudy (anonymous):

q1^2-q1(-7E-6)+(21.5E-11)=0 ?

OpenStudy (anonymous):

you're close but it's not quite what I have

OpenStudy (anonymous):

\[q1+q2=7*10^{-6}\]

OpenStudy (anonymous):

\[q1*q2=1.027*10^{-12}\]

OpenStudy (anonymous):

im getting lost in the algebra substitute q1 or 2 for a and solve right?

OpenStudy (anonymous):

try working through it without numbers first, so say q1 + q2 = m and q1*q2 = n

OpenStudy (anonymous):

so q2 = n/q1

OpenStudy (anonymous):

then substitute for q2 in the first expression to get q1 +n/q1 =m

OpenStudy (anonymous):

multiply through by q1 to get q1^2 +n = m*q1

OpenStudy (anonymous):

rearrange to get q1^2 - m*q1 + n = 0

OpenStudy (anonymous):

then the solutions are \[\frac{ 1 }{ 2}[m \pm \sqrt{m^2-4*n}]\]

OpenStudy (anonymous):

plug in the values, \[m=7*10^{-6}\] and \[n=1.027*10^{-12}\]

OpenStudy (anonymous):

let me know what you get for the two solutions

OpenStudy (anonymous):

6.82 and1.027

OpenStudy (anonymous):

well i have 6.85 for one value

OpenStudy (anonymous):

do your two values add up to 7 like they're supposed to ?

OpenStudy (anonymous):

i was * to make 7

OpenStudy (anonymous):

.18

OpenStudy (anonymous):

ok, if you try 6.82 and o.18, what do they give for the product q1*q2 ?

OpenStudy (anonymous):

when i enter my answers it tells me 6.82 is correct for the large charge but 0.18 is wrong for the smaller

OpenStudy (anonymous):

the correct answers are 6.85 and 0.15, you've probably made a slight error with the numbers somewhere

OpenStudy (anonymous):

microcoulombs, of course

OpenStudy (anonymous):

yes the the -6

OpenStudy (anonymous):

q1*q2 has to come to 1.027 *10-12 and q1+q2 has to come to 7*10-6

OpenStudy (anonymous):

it accepted .15 wth strange

OpenStudy (anonymous):

and the 6.82

OpenStudy (anonymous):

i thought they should have equaled 7 but

OpenStudy (anonymous):

the main thing is to understand the method, you might want to check your numerical calculation to see if there was a slight error somewhere

OpenStudy (anonymous):

yes they do have to equal 7, the larger charge is 6.85

OpenStudy (anonymous):

thanks for spending so much time with me

OpenStudy (anonymous):

no problem

OpenStudy (anonymous):

so now i do the same but with opp charge right

OpenStudy (anonymous):

what about the attractive force ?

OpenStudy (anonymous):

use the numbers i just got?

OpenStudy (anonymous):

the question states that the force is repulsive, you have the two values of the charges, there is no more to be done, you've solved the problem

OpenStudy (anonymous):

there is a part b asking the charge on force is attractive

OpenStudy (anonymous):

but thats cool i dont wanna bug ya i appreciate what youve helped me with